# The Fefferman metric for twistor CR manifolds and conformal geodesics in dimension three

### v3

@@ 2.1. CR structures
$T^{1,0}_{x}M\times T^{0,1}_{x}M\ni(Z,\overline{W})\longmapsto i[Z,\overline{W}]\in T_{x}M/H_{x}$
@@
We usually assume that the real line bundle $TM/H$ is orientable so that we can trivialize $H^{\perp}\subset T^{*}M$ by a global contact form $\theta\in\Gamma(H^{\perp})$. For a fixed orientation of $TM/H$, a contact form is determined up to multiplications by positive functions: $\theta\mapsto\widehat{\theta}=e^{\Upsilon}\theta,\Upsilon\in C^{\infty}(M)$. For each choice of $\theta$, we have the real valued Levi form $h_{\theta}$ by trivializing $TM/H$, which transforms conformally as $h_{\widehat{\theta}}=e^{\Upsilon}h_{\theta}$.
The Reeb vector field of a contact form $\theta$ is a real vector field $T$ uniquely specified by the conditions $\theta(T)=1,T\lrcorner\,d\theta=0$. Choosing a local frame $(Z_{\alpha})$ for $T^{1,0}M$, we have a local frame $(T,Z_{\alpha},Z_{\overline{\alpha}}:=\overline{Z_{\alpha}})$ for $\mathbb{C}TM=\mathbb{C}T\oplus T^{1,0}M\oplus T^{0,1}M$ called an admissible frame, and its dual coframe $(\theta,\theta^{\alpha},\theta^{\overline{\alpha}}=\overline{\theta^{\alpha}})$. In this coframe, we can write as
$d\theta=ih_{\alpha\overline{\beta}}\theta^{\alpha}\wedge\theta^{\overline{\beta}}$
and the Levi form is given by $(Z,\overline{W})\mapsto h_{\alpha\overline{\beta}}Z^{\alpha}\overline{W^{\beta}}$.
2.2. The Fefferman metric
@@
Let $(M,T^{1,0}M)$ be a non-degenerate $(2n+1)$-dimensional CR manifold. The CR canonical bundle is the complex line bundle over $M$ defined by
$K_{M}:=\Bigl{\{}\xi\in\bigwedge\nolimits^{\!n+1}(\mathbb{C}T^{*}M)\,\Big{|}\,\overline{Z}\lrcorner\,\xi=0\ \textrm{for any}\ \overline{Z}\in T^{0,1}M\Bigr{\}}.$
@@
(2.1) $\mathcal{C}\cong\{e^{is}\theta\wedge\theta^{1}\wedge\cdots\wedge\theta^{n}\,|\,s\in\mathbb{R}/2\pi\mathbb{Z}\}\subset K^{\times}_{M},$
where $(\theta,\theta^{\alpha},\theta^{\overline{\alpha}})$ is an admissible coframe such that $|\det(h_{\alpha\overline{\beta}})|=1$. This realization defines a tautological $(n+1)$-form
$\xi=e^{is}\theta\wedge\theta^{1}\wedge\cdots\wedge\theta^{n}$
on $\mathcal{C}$. Also, $\varphi=e^{is}\theta^{1}\wedge\cdots\wedge\theta^{n}$ becomes a global $n$-form on $\mathcal{C}$ uniquely characterized by the conditions
$\xi=\theta\wedge\varphi,\quad V\lrcorner\,\varphi=0\ \mathrm{for\ any\ lift}\ V\ \mathrm{of}\ T;$
see [18, Lemma 3.3].
We also have a $1$-form $\sigma$ on $\mathcal{C}$ specified by the following proposition:
Proposition 2.1 ([18, Proposition 3.4]).
@@
$d\xi$ $=i(n+2)\sigma\wedge\xi,$
$\sigma\wedge d\varphi\wedge\overline{\varphi}$ $=(\mathrm{Tr}\,d\sigma)i\sigma\wedge\theta\wedge\varphi\wedge\overline{\varphi}.$
Here, for a 2-form $\omega$ on $M$ written as $\omega=i\omega_{\alpha\overline{\beta}}\theta^{\alpha}\wedge\theta^{\overline{\beta}}+\cdots$ in an admissible coframe, we define ${\rm Tr}\,\omega:=h^{\alpha\overline{\beta}}\omega_{\alpha\overline{\beta}}$ using the inverse $h^{\alpha\overline{\beta}}$ of the Levi form. The $1$-form $(n+2)\sigma$ gives a principal connection form on the $S^{1}$-bundle $\mathcal{C}$.
@@
We note that our definition of $g^{\mathrm{F}}$ is twice the definition of Lee. In the coframe $(\theta^{0}:=\theta,\theta^{\alpha},\theta^{\overline{\alpha}},\theta^{n+1}:=2\sigma)$, the non-zero components of the metric tensor are
[…]
If the Levi form has signature $(p,q)$ as a hermitian form, then the Fefferman metric has signature $(2p+1,2q+1)$. Moreover, if we rescale the contact form as $\widehat{\theta}=e^{\Upsilon}\theta$, the Fefferman metric transforms as $\widehat{g}^{\,\mathrm{F}}=e^{\Upsilon}g^{\mathrm{F}}$. Thus, the conformal structure $[g^{\mathrm{F}}]$ is independent of the choice of contact form ([18, Theorem 5.17]).
@@
$\mathcal{Q}=\bigl{\{}[\xi]\in\mathbb{CP}^{n+1}\,\big{|}\,-|\xi^{0}|^{2}-|\xi^{1}|^{2}-\cdots-|\xi^{q}|^{2}+|\xi^{q+1}|^{2}+\cdots+|\xi^{n+1}|^{2}=0\bigr{\}},$
@@
2.3. Chains and null chains
@@
An unparametrized curve on $M$ is called a chain (resp. null chain) if it is the projection of a non-vertical null geodesic $\gamma(t)$ of $[g^{\mathrm{F}}]$ with $g^{\mathrm{F}}(K,\dot{\gamma})\neq 0$ (resp. $g^{\mathrm{F}}(K,\dot{\gamma})=0$).
@@
For each choice of a direction transverse to $H_{x}$, there exists a unique chain emanating in that direction. Thus, chains can be considered as ‘geodesics’ in CR geometry. In fact, they are characterized as geodesics of a Kropina metric on CR manifolds, which is a Finsler metric singular along $H$; see [6] and §5.2 below. On the other hand, there exists a 1-parameter family of null chains which are tangent to a fixed null vector in $H_{x}$; see [13].
3. Conformal three-manifolds and twistor CR manifolds
3.1. Conformal three-manifolds
@@
The Cotton tensor $C=(C_{ijk})\in\Gamma(T^{*}\Sigma\otimes\wedge^{2}T^{*}\Sigma)$ is defined by
$C(X,Y,Z):=(\nabla_{Z}P)(X,Y)-(\nabla_{Y}P)(X,Z),\ \textrm{i.e.,}\ C_{ijk}=\nabla_{k}P_{ij}-\nabla_{j}P_{ik}.$
Since $\Sigma$ is 3-dimensional, the Cotton tensor is conforamlly invariant, and identically vanishes if and only if $(\Sigma,[g])$ is conformally flat.
@@
Applying the Hodge star operator to the Cotton tensor, we define $*C\in\Gamma(T^{*}\Sigma\otimes T^{*}\Sigma)$ by
@@ 3.2. Conformal geodesics
(3.2) $\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-\frac{3g(\dot{x},\nabla_{\dot{x}}\dot{x})}{|\dot{x}|^{2}}\nabla_{\dot{x}}\dot{x}+\frac{3|\nabla_{\dot{x}}\dot{x}|^{2}}{2|\dot{x}|^{2}}\dot{x}+2P(\dot{x},\dot{x})\dot{x}-|\dot{x}|^{2}P(\dot{x})=0.$
@@
It is shown in [1] that if $x(t)$ is a solution to the equation (3.2), a reparametrized curve $\tilde{x}(t)=x(\varphi(t))$ solves (3.2) if and only if $\varphi(t)$ is a projective linear transformation: $\varphi(t)=(at+b)/(ct+d)$. Hence, we call a solution to (3.2) a conformal geodesic with a projective parameter. In [1], they also prove that one can define an equation for ‘unparametrized conformal geodesics’ by taking the $\dot{x}$-orthogonal part of (3.2):
@@
(3.3) $\dot{x}\wedge\left(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-\frac{3g(\dot{x},\nabla_{\dot{x}}\dot{x})}{|\dot{x}|^{2}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x})\right)=0.$
@@
(3.4) $\dot{x}\wedge\left(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x})\right)=0$
@@
Conformal geodesics are also characterized as follows:
@@
A regular curve $x(t)$ is a conformal geodesic if and only if there exists a metric $g\in[g]$ and a reparametrization such that
@@
3.3. Twistor CR manifolds
We will define the twistor CR manifold $M$ over $(\Sigma^{3},[g])$, which was introduced by LeBrun [17] as a $3$-dimensional analogue of Penrose’s twistor space for anti-self dual conformal $4$-manifolds.
@@ 3.3. Twistor CR manifolds
$\hat{M}:=\{\zeta\in\mathbb{C}T^{*}\Sigma\ |\ g(\zeta,\zeta)=0,\zeta\neq 0\}.$
By taking the quotient by the $\mathbb{C}^{*}$-action, we obtain a $5$-dimentional manifold
@@
We denote the natural projections by $\tilde{\pi}\colon\hat{M}\to M,\ \pi\colon M\to\Sigma$. Note that the fiber of $\pi$ is biholomorphic to $\{[\zeta]\in\mathbb{CP}^{2}\,|\,\zeta_{1}^{2}+\zeta_{2}^{2}+\zeta_{3}^{2}=0\}\cong\mathbb{CP}^{1}$.
Let $\alpha:=(\zeta_{i}dx^{i})|_{T\hat{M}}$ be the canonical (tautological) $1$-form on $\hat{M}$ and set $\omega:=d\alpha$. Then, the distribution
$D:=\textrm{Ker}\,\omega=\{v\in\mathbb{C}T\hat{M}\ |\ v\lrcorner\,\omega=0\}$
@@
becomes a non-degenerate CR structure on $M$ whose Levi form has signature $(1,1)$. The construction depends only on the conformal class $[g]$, and $(M,T^{1,0}M)$ is called the twistor CR manifold over $(\Sigma,[g])$.
@@
For any point $[\zeta]\in\pi^{-1}(x_{0})$, we have
$\overline{D}_{\zeta}=\mathbb{C}\overline{\zeta^{i}}\frac{\partial}{\partial x^{i}}\oplus\Bigl{\{}w_{i}\frac{\partial}{\partial\zeta_{i}}\,\Big{|}\,\zeta_{i}w^{i}=0\Bigr{\}},$
@@
$T^{1.0}_{[\zeta]}M\cong\mathbb{C}\overline{\zeta^{i}}\frac{\partial}{\partial x^{i}}\oplus\left(\Bigl{\{}w_{i}\frac{\partial}{\partial\zeta_{i}}\,\Big{|}\,\zeta_{i}w^{i}=0\Bigr{\}}\Big{/}\mathbb{C}\zeta_{i}\frac{\partial}{\partial\zeta_{i}}\right).$
See [20, Proposition 6.1] for the proof.
4. The Fefferman metric and (null) chains of twistor CR manifolds
@@
$\widehat{\mathcal{C}}:=\{\zeta\in\hat{M}\ |\ g(\zeta,\overline{\zeta})=1\}=\{\zeta\in\mathbb{C}T^{*}\Sigma\ |\ g(\zeta,\zeta)=0,\,g(\zeta,\overline{\zeta})=1\},$
@@
4.1. An adapted frame on $\widehat{\mathcal{C}}$
Let $(\zeta_{i})$ be the fiber coordinates for $\zeta_{i}dx^{i}\in\mathbb{C}T^{*}\Sigma$ associated to local coordinates $(x^{i})$ on $\Sigma$. On $\widehat{\mathcal{C}}$, it holds that
$\zeta^{i}\zeta_{i}=0,\quad\zeta^{i}\overline{\zeta}_{i}=1.$
By using the metric, we identify $\zeta_{i}dx^{i}$ with the complex tangent vector

### v4

@@ 2.1 CR structures
$T^{1,0}_{x}M\times T^{0,1}_{x}M\ni\bigl(Z,\overline{W}\bigr)\longmapsto{\rm i}\bigl[Z,\overline{W}\bigr]\in T_{x}M/H_{x}$
@@
We usually assume that the real line bundle $TM/H$ is orientable so that we can trivialize $H^{\perp}\subset T^{*}M$ by a global contact form $\theta\in\Gamma\bigl(H^{\perp}\bigr)$. For a fixed orientation of $TM/H$, a contact form is determined up to multiplications by positive functions: $\theta\mapsto\widehat{\theta}={\rm e}^{\Upsilon}\theta$, $\Upsilon\in C^{\infty}(M)$. For each choice of $\theta$, we have the real valued Levi form $h_{\theta}$ by trivializing $TM/H$, which transforms conformally as $h_{\widehat{\theta}}={\rm e}^{\Upsilon}h_{\theta}$.
The Reeb vector field of a contact form $\theta$ is a real vector field $T$ uniquely specified by the conditions $\theta(T)=1$, $T\lrcorner\,{\rm d}\theta=0$. Choosing a local frame $(Z_{\alpha})$ for $T^{1,0}M$, we have a local frame $\bigl(T,Z_{\alpha},Z_{\overline{\alpha}}:=\overline{Z_{\alpha}}\bigr)$ for $\mathbb{C}TM=\mathbb{C}T\oplus T^{1,0}M\oplus T^{0,1}M$ called an admissible frame, and its dual coframe $\bigl(\theta,\theta^{\alpha},\theta^{\overline{\alpha}}=\overline{\theta^{\alpha}}\bigr)$. In this coframe, we can write as
${\rm d}\theta={\rm i}h_{\alpha\overline{\beta}}\theta^{\alpha}\wedge\theta^{\overline{\beta}}$
and the Levi form is given by $\bigl(Z,\overline{W}\bigr)\mapsto h_{\alpha\overline{\beta}}Z^{\alpha}\overline{W^{\beta}}$.
2.2 The Fefferman metric
@@
Let $\bigl(M,T^{1,0}M\bigr)$ be a non-degenerate $(2n+1)$-dimensional CR manifold. The CR canonical bundle is the complex line bundle over $M$ defined by
$K_{M}:=\biggl\{\xi\in\bigwedge\nolimits^{\!n+1}(\mathbb{C}T^{*}M)\,\Big|\,\overline{Z}\lrcorner\,\xi=0\ \textrm{for any}\ \overline{Z}\in T^{0,1}M\biggr\}.$
@@
$\mathcal{C}\cong\bigl\{{\rm e}^{{\rm i}s}\theta\wedge\theta^{1}\wedge\cdots\wedge\theta^{n}\mid s\in\mathbb{R}/2\pi\mathbb{Z}\bigr\}\subset K^{\times}_{M},$ (2.1)
where $\bigl(\theta,\theta^{\alpha},\theta^{\overline{\alpha}}\bigr)$ is an admissible coframe such that $\bigl|\det\bigl(h_{\alpha\overline{\beta}}\bigr)\bigr|=1$. This realization defines a tautological $(n+1)$-form
$\xi={\rm e}^{{\rm i}s}\theta\wedge\theta^{1}\wedge\cdots\wedge\theta^{n}$
on $\mathcal{C}$. Also, $\varphi={\rm e}^{{\rm i}s}\theta^{1}\wedge\cdots\wedge\theta^{n}$ becomes a global $n$-form on $\mathcal{C}$ uniquely characterized by the conditions
$\xi=\theta\wedge\varphi,\qquad V\lrcorner\,\varphi=0\quad\text{for any lift}\ V\ \text{of}\ T;$
see [22, Lemma 3.3].
We also have a $1$-form $\sigma$ on $\mathcal{C}$ specified by the following proposition.
Proposition 2.1 ([22, Proposition 3.4]).
@@
${\rm d}\xi={\rm i}(n+2)\sigma\wedge\xi,$
$\sigma\wedge{\rm d}\varphi\wedge\overline{\varphi}=(\operatorname{Tr}{\rm d}\sigma){\rm i}\sigma\wedge\theta\wedge\varphi\wedge\overline{\varphi}.$
Here, for a 2-form $\omega$ on $M$ written as $\omega={\rm i}\omega_{\alpha\overline{\beta}}\theta^{\alpha}\wedge\theta^{\overline{\beta}}+\cdots$ in an admissible coframe, we define $\operatorname{Tr}\omega:=h^{\alpha\overline{\beta}}\omega_{\alpha\overline{\beta}}$ using the inverse $h^{\alpha\overline{\beta}}$ of the Levi form. The $1$-form $(n+2)\sigma$ gives a principal connection form on the $S^{1}$-bundle $\mathcal{C}$.
@@
We note that our definition of $g^{\mathrm{F}}$ is twice the definition of Lee. In the coframe $\smash{\bigl(\theta^{0}:=\theta,\theta^{\alpha},\theta^{\overline{\alpha}}},\allowbreak\smash{\theta^{n+1}:=2\sigma\bigr)}$, the non-zero components of the metric tensor are
$g^{\mathrm{F}}_{\alpha\overline{\beta}}=g^{\mathrm{F}}_{\overline{\beta}\alpha}=h_{\alpha\overline{\beta}},\qquad g^{\mathrm{F}}_{0\,n+1}=g^{\mathrm{F}}_{n+1\,0}=1.$
If the Levi form has signature $(p,q)$ as a hermitian form, then the Fefferman metric has signature $(2p+1,2q+1)$. Moreover, if we rescale the contact form as $\widehat{\theta}={\rm e}^{\Upsilon}\theta$, the Fefferman metric transforms as $\widehat{g}^{\,\mathrm{F}}={\rm e}^{\Upsilon}g^{\mathrm{F}}$. Thus, the conformal structure $\bigl[g^{\mathrm{F}}\bigr]$ is independent of the choice of contact form [22, Theorem 5.17].
@@
$\mathcal{Q}=\bigl\{[\xi]\in\mathbb{CP}^{n+1}\,\big|\,-\bigl|\xi^{0}\bigr|^{2}-\bigl|\xi^{1}\bigr|^{2}-\cdots-\bigl|\xi^{q}\bigr|^{2}+\bigl|\xi^{q+1}\bigr|^{2}+\cdots+\bigl|\xi^{n+1}\bigr|^{2}=0\bigr\},$
@@
2.3 Chains and null chains
@@
An unparametrized curve on $M$ is called a chain (resp. null chain) if it is the projection of a non-vertical null geodesic $\gamma(t)$ of $\bigl[g^{\mathrm{F}}\bigr]$ with $g^{\mathrm{F}}(K,\dot{\gamma})\neq 0$ (resp. $g^{\mathrm{F}}(K,\dot{\gamma})=0$).
@@
For each choice of a direction transverse to $H_{x}$, there exists a unique chain emanating in that direction. Thus, chains can be considered as ‘geodesics’ in CR geometry. In fact, they are characterized as geodesics of a Kropina metric on CR manifolds, which is a Finsler metric singular along $H$; see [9] and Section 5.2 below. On the other hand, there exists a 1-parameter family of null chains which are tangent to a fixed null vector in $H_{x}$; see [17].
3 Conformal three-manifolds and twistor CR manifolds
3.1 Conformal three-manifolds
@@
The Cotton tensor $C=(C_{ijk})\in\Gamma\bigl(T^{*}\Sigma\otimes\wedge^{2}T^{*}\Sigma\bigr)$ is defined by
$C(X,Y,Z):=(\nabla_{Z}P)(X,Y)-(\nabla_{Y}P)(X,Z),\qquad\text{i.e.,}\qquad C_{ijk}=\nabla_{k}P_{ij}-\nabla_{j}P_{ik}.$
Since $\Sigma$ is 3-dimensional, the Cotton tensor is conformally invariant, and identically vanishes if and only if $(\Sigma,[g])$ is conformally flat.
@@
Applying the Hodge star operator to the Cotton tensor, we define $C\in\Gamma(T^{*}\Sigma\otimes T^{*}\Sigma)$ by
@@ 3.2 Conformal geodesics
$\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-\frac{3g(\dot{x},\nabla_{\dot{x}}\dot{x})}{|\dot{x}|^{2}}\nabla_{\dot{x}}\dot{x}+\frac{3|\nabla_{\dot{x}}\dot{x}|^{2}}{2|\dot{x}|^{2}}\dot{x}+2P(\dot{x},\dot{x})\dot{x}-|\dot{x}|^{2}P(\dot{x})=0.$ (3.2)
@@
It is shown in [1] that if $x(t)$ is a solution to the equation (3.2), a reparametrized curve $\tilde{x}(t)=x(\varphi(t))$ solves (3.2) if and only if $\varphi(t)$ is a projective linear transformation: $\varphi(t)=(at+b)/(ct+d)$. Hence, we call a solution to (3.2) a conformal geodesic with a projective parameter. In [1], they also prove that one can define an equation for ‘unparametrized conformal geodesics’ by taking the $\dot{x}$-orthogonal part of (3.2).
@@
$\dot{x}\wedge\biggl(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-\frac{3g(\dot{x},\nabla_{\dot{x}}\dot{x})}{|\dot{x}|^{2}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x})\biggr)=0.$ (3.3)
@@
$\dot{x}\wedge\bigl(\nabla_{\dot{x}}\nabla_{\dot{x}}\dot{x}-|\dot{x}|^{2}P(\dot{x})\bigr)=0$ (3.4)
@@
Conformal geodesics are also characterized as follows.
@@
A regular curve $x(t)$ is a conformal geodesic if and only if there exist a metric $g\in[g]$ and a reparametrization such that
@@
3.3 Twistor CR manifolds
We will define the twistor CR manifold $M$ over $\bigl(\Sigma^{3},[g]\bigr)$, which was introduced by LeBrun [21] as a $3$-dimensional analogue of Penrose’s twistor space for anti-self dual conformal $4$-manifolds.
@@ 3.3 Twistor CR manifolds
$\hat{M}:=\{\zeta\in\mathbb{C}T^{*}\Sigma\mid g(\zeta,\zeta)=0,\,\zeta\neq 0\}.$
By taking the quotient by the $\mathbb{C}^{*}$-action, we obtain a $5$-dimensional manifold
@@
We denote the natural projections by $\tilde{\pi}\colon\hat{M}\to M$, $\pi\colon M\to\Sigma$. Note that the fiber of $\pi$ is biholomorphic to $\bigl\{[\zeta]\in\mathbb{CP}^{2}\mid\zeta_{1}^{2}+\zeta_{2}^{2}+\zeta_{3}^{2}=0\bigr\}\cong\mathbb{CP}^{1}$.
Let $\alpha:=\bigl(\zeta_{i}{\rm d}x^{i}\bigr)|_{T\hat{M}}$ be the canonical (tautological) $1$-form on $\hat{M}$ and set $\omega:={\rm d}\alpha$. Then, the distribution
$D:=\textrm{Ker}\,\omega=\bigl\{v\in\mathbb{C}T\hat{M}\mid v\lrcorner\,\omega=0\bigr\}$
@@
becomes a non-degenerate CR structure on $M$ whose Levi form has signature $(1,1)$. The construction depends only on the conformal class $[g]$, and $\bigl(M,T^{1,0}M\bigr)$ is called the twistor CR manifold over $(\Sigma,[g])$.
@@
For any point $\zeta]\in\pi^{-1}(x_{0})$, we have
$\overline{D}_{\zeta}=\mathbb{C}\overline{\zeta^{i}}\frac{\partial}{\partial x^{i}}\oplus\biggl\{w_{i}\frac{\partial}{\partial\zeta_{i}}\,\bigg|\,\zeta_{i}w^{i}=0\biggr\},$
@@
$T^{1,0}_{[\zeta]}M\cong\mathbb{C}\overline{\zeta^{i}}\frac{\partial}{\partial x^{i}}\oplus\biggl(\biggl\{w_{i}\frac{\partial}{\partial\zeta_{i}}\,\bigg|\,\zeta_{i}w^{i}=0\biggr\}\Big/\mathbb{C}\zeta_{i}\frac{\partial}{\partial\zeta_{i}}\biggr).$
See [24, Proposition 6.1] for the proof.
4 The Fefferman metric and (null) chains
of twistor CR manifolds
@@ 4 The Fefferman metric and (null) chains
$\widehat{\mathcal{C}}:=\bigl\{\zeta\in\hat{M}\mid g\bigl(\zeta,\overline{\zeta}\bigr)=1\bigr\}=\bigl\{\zeta\in\mathbb{C}T^{*}\Sigma\mid g(\zeta,\zeta)=0,\,g\bigl(\zeta,\overline{\zeta}\bigr)=1\bigr\},$
@@
4.1 An adapted frame on $\boldsymbol{\widehat{\mathcal{C}}}$
Let $(\zeta_{i})$ be the fiber coordinates for $\zeta_{i}{\rm d}x^{i}\in\mathbb{C}T^{*}\Sigma$ associated to local coordinates $(x^{i})$ on $\Sigma$. On $\widehat{\mathcal{C}}$, it holds that
$\zeta^{i}\zeta_{i}=0,\qquad\zeta^{i}\overline{\zeta}_{i}=1.$
By using the metric, we identify $\zeta_{i}{\rm d}x^{i}$ with the complex tangent vector